OpenStudy (anonymous):

Which one is a better buy? 10 candy bars for $6.75 or 12 candy bars for $7.25? please help i'll medal and fan. @bohotness

OpenStudy (anonymous):

ok so what is 10/6.75?

OpenStudy (anonymous):

that is how much 1 candy bar costs

OpenStudy (perl):

compare unit rates 6.75 / 10 and 7.25 / 12

OpenStudy (anonymous):

then you do 12/7.25

OpenStudy (anonymous):

then compare them and you will figure out the better buy

OpenStudy (anonymous):

I don't get it.

OpenStudy (perl):

make sure the units are correct you want dollars / item

OpenStudy (briellshaw):

I would think 12 candy bars for $7.25 because it a cheap offer that has more usually 12 candy bars is over $10.00

OpenStudy (anonymous):

so 10/6.75=1.49

OpenStudy (perl):

i cant use the draw button :/

OpenStudy (anonymous):

and 12/7.25=1.66

OpenStudy (anonymous):

so which is the better buy

OpenStudy (anonymous):

oh ok so it would be the 10 for 6.75 right?

OpenStudy (anonymous):

sorry i might of done that wrong try this: 6/27=2/9 2*3=6 9*3=27 6/27 21. 6.75/10=.68c (c=cent(s) 7.25/12=61c

OpenStudy (perl):